OpenStudy (anonymous):

Simplify completely quantity 6 x squared minus 54 x plus 84 over quantity 8 x squared minus 40 x plus 48 divided by quantity x squared plus x minus 56 over quantity 2 x squared plus 12 x minus 32

OpenStudy (anonymous):

\[\frac{ 6x^2-54x+84 }{8x^2-40x+48 }\div \frac{ x^2+x-56 }{ 2x^2+12x-32 }\]

OpenStudy (anonymous):

type me what you worked out so far

OpenStudy (anonymous):

\[\frac{6(x-7)(x-2) }{8(x-6)(x+1) }\times \frac{ 2(x+8)(x-2) }{ (x+8)(x-7) }\]

OpenStudy (anonymous):

,flip the frations on the right side

OpenStudy (anonymous):

*fraction

OpenStudy (anonymous):

What does that change? Besides the numerator and denominator being in different positions?

OpenStudy (anonymous):

flip the fraction on the right, the you can cancel out, then multiply

OpenStudy (anonymous):

Do I cancel out the two x-2's?

OpenStudy (anonymous):

damn, i gtg, message me when i ge back and ill walk you through this

OpenStudy (anonymous):

Okay thanks